OpenStudy (anonymous):

Line AB contains points A (8, āˆ’4) and B (1, āˆ’5). The slope of line AB is... āˆ’7 negative 1 over 7 1 over 7 7

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&({\color{red}{ 8}}\quad ,&{\color{blue}{ -4}})\quad B&({\color{red}{ 1}}\quad ,&{\color{blue}{ -5}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (anonymous):

all i have to do is the y2 -y1 thingy??? @jdoe0001

OpenStudy (jdoe0001):

yes, to find the slope of a line from 2 points

OpenStudy (anonymous):

is it b???@jdoe0001

OpenStudy (jdoe0001):

well.. what did you get from \(\bf \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\quad ?\)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&({\color{red}{ 8}}\quad ,&{\color{blue}{ -4}})\quad B&({\color{red}{ 1}}\quad ,&{\color{blue}{ -5}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ -5}}-{\color{blue}{ (-4)}}}{{\color{red}{ 1}}-{\color{red}{ 8}}}\implies ?\)

OpenStudy (anonymous):

i got b @jdoe0001

OpenStudy (anonymous):

am i correct??@jdoe0001

OpenStudy (jdoe0001):

hmm what did you get for the numerator?

OpenStudy (anonymous):

numerator is top right?? @jdoe0001

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

okay i got -1.. @jdoe0001

OpenStudy (anonymous):

-1/7

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&({\color{red}{ 8}}\quad ,&{\color{blue}{ -4}})\quad B&({\color{red}{ 1}}\quad ,&{\color{blue}{ -5}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ -5}}-{\color{blue}{ (-4)}}}{{\color{red}{ 1}}-{\color{red}{ 8}}}\implies \cfrac{-5+4}{-7}\implies \cfrac{-1}{-7} \\ \quad \\ \implies \cfrac{1}{7}\)

OpenStudy (anonymous):

ohhh okay yeahh i got it because a - and a - is a positive thanks!!! @jdoe0001

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

yw